### v2

# A Comprehensive Survey on Self-Supervised Learning for Recommendation

Abstract.
Recommender systems play a crucial role in tackling the challenge of information overload by delivering personalized recommendations based on individual user preferences. Deep learning techniques, such as RNNs, GNNs, and Transformer architectures, have significantly propelled the advancement of recommender systems by enhancing their comprehension of user behaviors and preferences. However, supervised learning methods encounter challenges in real-life scenarios due to data sparsity, resulting in limitations in their ability to learn representations effectively. To address this, self-supervised learning (SSL) techniques have emerged as a solution, leveraging inherent data structures to generate supervision signals without relying solely on labeled data. By leveraging unlabeled data and extracting meaningful representations, recommender systems utilizing SSL can make accurate predictions and recommendations even when confronted with data sparsity. In this paper, we provide a comprehensive review of self-supervised learning frameworks designed for recommender systems, encompassing a thorough analysis of over 170 papers. We conduct an exploration of nine distinct scenarios, enabling a comprehensive understanding of SSL-enhanced recommenders in different contexts. For each domain, we elaborate on different self-supervised learning paradigms, namely contrastive learning, generative learning, and adversarial learning, so as to present technical details of how SSL enhances recommender systems in various contexts. We consistently maintain the related open-source materials at https://github.com/HKUDS/Awesome-SSLRec-Papers.
Recommendation, Self-Supervised Learning
copyright: acmcopyrightccs: General and reference Surveys and overviewsccs: Information systems Recommender systems
1. Introduction
Recommender systems play a vital role in addressing the challenge of information overload by providing personalized recommendations to individual users based on their unique preferences (Zhang et al., 2019). These systems are designed to enhance the overall user experience by presenting users with recommendations that are not only relevant but also closely aligned with their interests. This tailored approach makes the user experience more engaging, efficient, and ultimately more satisfying. At the core of recommender systems lies the fundamental principle of understanding users’ preferences for a diverse range of items (Mao et al., 2021; Lai et al., 2023). This understanding is achieved through a meticulous analysis of users’ past interactions, which encompass activities such as clicks and purchases. By examining these interactions, recommender systems gain valuable insights into users’ behavior, enabling them to identify patterns and uncover individual preferences.
The field of recommender systems has undergone a revolutionary transformation, thanks to the strong representation capabilities of deep learning techniques. Neural network architectures have played a pivotal role in this paradigm shift. By harnessing the power of deep learning models, such as Recurrent Neural Networks (RNNs) (Jannach and Ludewig, 2017; Guo et al., 2020), Graph Neural Networks (GNNs) (Ying et al., 2018; Wang et al., 2019a), and Transformer architectures (Sun et al., 2019; Wu et al., 2020a), recommender systems have achieved an unprecedented level of understanding of user preferences and behaviors. Consequently, this deep understanding has paved the way for precise and personalized recommendations that cater to individual user needs and preferences.
Existing supervised learning methods heavily depend on having abundant labeled data for effective training. However, practical recommender systems often encounter the problem of data sparsity (Xia et al., 2022b). This means that real-life recommendation scenarios often suffer from a limited amount of available data or an insufficient number of labeled examples. Consequently, supervised learning methods encounter significant difficulties in effectively generalizing and accurately learning representations of user preferences in such scenarios. Fortunately, taking inspiration from the accomplishments of self-supervised learning (SSL) (Liu et al., 2021d), SSL techniques has proven to be beneficial for recommender systems in addressing the issue of data sparsity (Ren et al., 2023b; Xia et al., 2022b). Generally speaking, the key idea of SSL lies in leveraging the inherent structures or patterns within the data itself to create supervision signals for learning, without relying solely on externally labeled data. This allows recommender systems to utilize unlabeled data and extract meaningful representations, enabling them to make accurate predictions and recommendations even in the presence of data sparsity.
Our paper provides a comprehensive review of the latest advancements in self-supervised learning frameworks tailored specifically for recommendation systems. It aims to serve as a valuable resource for researchers from diverse disciplines, extending beyond the realms of computer science and machine learning, who wish to explore this rapidly evolving field. In this context, our paper presents several significant contributions, which are summarized as follows:
Comprehensive Collection of Papers. We have conducted a thorough review of over 170 papers that investigate the application of self-supervised learning in the field of recommendation. Our search was carried out using reputable academic databases such as Google Scholar and DBLP, utilizing specific keywords including “self-supervised”, “contrastive”, “generative”, “adversarial”, “variational”, “diffusion” and “masked autoencoder” in conjunction with “recommendation” and “recommender systems”. The surveyed papers were meticulously sourced from esteemed conferences and journals such as KDD, SIGIR, WWW, ICLR, WSDM, CIKM, ICDE, AAAI, IJCAI, RecSys, TOIS, TKDE. To ensure the inclusion of state-of-the-art research, we also explored citation networks and incorporated relevant preprints from arXiv.
Supported with Open-source Library. Our team has developed SSLRec (Ren et al., 2024b), a robust framework for self-supervised learning. This user-friendly framework encompasses popular datasets, standardized code scripts for data processing, training, testing, evaluation, and cutting-edge recommender models for self-supervised learning.
Relationship with Previous Surveys. i) Extensive Collection of SSL Research. Our survey provides a significantly broader collection of self-supervised learning (SSL) works in the context of recommendation, encompassing over 170 papers, surpassing the previous surveys which covered approximately 80 papers (Yu et al., 2023b; Jing et al., 2023). In addition, we incorporate generative methods (e.g., mask autoencoding, denoised diffusion) and adversarial learning, which were not previously covered but have gained prominence according to (Liu et al., 2021d). ii) Comprehensive Taxonomy Design. Our taxonomy refines previous classifications of contrastive learning (Yu et al., 2023; Jing et al., 2023) and proposes a view-centric taxonomy approach. We also survey generative and adversarial learning methods based on their learning paradigms and targets. iii) Exploration of Distinct Scenarios. Distinguishing ourselves from earlier works (Yu et al., 2023b; Jing et al., 2023), which did not differentiate SSL-enhanced recommenders across different scenarios, we conducted an extensive exploration of nine scenarios individually. By surveying research works within diverse scenarios, we provide researchers with a more comprehensive understanding of the context-specific to each scenario and the corresponding challenges they present.
Organization of the Survey: The structure of the survey is as follows: Section 2 provides a comprehensive overview of recommendation systems and self-supervised learning, establishing the necessary background knowledge. In Section 3, we present our proposed taxonomy for understanding self-supervised learning within the context of recommendation systems. The main content of the survey is covered in Section 4, where we conduct separate reviews of the three self-supervised learning paradigms across various recommendation scenarios. Moving forward, Section 5 delves into open problems and future directions in this field. Lastly, we conclude the survey with Section 6.
2. Preliminaries and Definition
In this section, we provide a concise introduction to the essential background knowledge relevant to our survey on self-supervised for learning recommendation. We begin by briefly outlining the preliminaries of the tasks in recommender systems. Subsequently, we describe the definition of self-supervised learning and introduce prominent learning paradigms: contrastive, generative, and adversarial approaches.
2.1. Recommender Systems
The research on recommendation encompasses a wide range of tasks in diverse scenarios, such as collaborative filtering, sequential recommendation, and multi-behavior recommendation. These tasks exhibit distinct data paradigms and objectives. Here, we begin by providing a general definition without delving into specific variations across diverse recommendation tasks. In recommender systems, there are two primary sets: the set of users, denoted as $\mathcal{U}=\{u_{1},...,u_{i},...,u_{|\mathcal{U}|}\}$, and the set of items, denoted as $\mathcal{V}=\{v_{1},...,v_{j},...,v_{|\mathcal{V}|}\}$. Then, an interaction matrix $\mathcal{A}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{V}|}$ is utilized to represent the recorded interactions between users and items. In this matrix, an entry $\mathcal{A}_{i,j}$ is assigned a value of $1$ if user $u_{i}$ has interacted with item $v_{j}$, and 00 otherwise. The definition of interaction is adaptable to various scenarios and datasets (e.g., watching a movie, clicking or making purchases on an e-commerce website). Furthermore, in various recommendation tasks, there are distinct auxiliary observed data denoted as $\mathcal{X}$. For instance, in knowledge graph-enhanced recommendation, $\mathcal{X}$ incorporates a knowledge graph comprising external item attributes with different entity types and corresponding relationships. While in social recommendation, $\mathcal{X}$ encompasses user-level relationships, such as friendships. With the above definition, a recommendation model optimizes a prediction function $f(\cdot)$ that aims to accurately estimate the preference score between any user $u$ and item $v$:
(1) $y_{u,v}=f(\mathcal{A},\mathcal{X},u,v).$
The preference score $y_{u,v}$ represents the likelihood of user $u$ interacting with item $v$. Based on this score, the recommender system can provide recommendations of uninteracted items to each user by presenting a ranking list of items according to the estimated preference scores. In Section 4, we will delve into the data formulation of $(\mathcal{A},\mathcal{X})$ under different recommendation scenarios, complementing the demonstration of self-supervised learning in recommendation.
2.2. Self-supervised Learning in Recommendation
Over the past years, deep neural networks have demonstrated outstanding performance with supervised learning in various fields, including computer vision (Voulodimos et al., 2018), natural language processing (Deng and Liu, 2018) and recommender systems (Zhang et al., 2019). However, due to the heavy reliance on labeled data, supervised learning faces challenges when dealing with label sparsity, which is also a common issue in recommendation systems (Wu et al., 2021b; Yu et al., 2023b). To address this limitation, self-supervised learning has emerged as a promising approach, leveraging the data itself as labels for learning. In this section, we introduce three fundamental self-supervised learning methodologies: contrastive, generative, and adversarial.
2.2.1. Contrastive Learning
To fully leverage the inherent information within the data itself as supervision signals, contrastive learning has emerged as a prominent self-supervised learning approach (Jaiswal et al., 2020). The primary objective of contrastive learning is to maximize the agreement between different views augmented from the data. Formally, in contrastive learning for the recommendation, the objective is to minimize the following loss function (Wu et al., 2021a; Xie et al., 2022c):
(2) $\mathop{\min}_{\mathcal{E}_{1},\mathcal{E}_{2}}\ \mathcal{L}_{con}(\mathcal{E}_{1}\circ\omega_{1}(\mathcal{A},\mathcal{X}),\ \mathcal{E}_{2}\circ\omega_{2}(\mathcal{A},\mathcal{X})).$
Here *$\mathcal{E}_{*}\circ\omega_{*}$ represents the view creation operations, which can vary across different contrastive learning-based recommenders. Each operation consists of data permutation processes, $\omega_{1}(\cdot)$ and $\omega_{2}(\cdot)$, which may involve dropping nodes/edges in graphs, as well as embedding encoding processes, $\mathcal{E}_{1}$ and $\mathcal{E}2$. The objective of minimizing $\mathcal{L}_{con}$ is to obtain robust encoding functions that maximize the agreement between the views. This cross-view agreement can be achieved through methods such as mutual information maximization or instance discrimination.
2.2.2. Generative Learning
Generative learning seeks to understand data structures and patterns to learn meaningful representations. It optimizes a deep encoder-decoder model that reconstructs missing or corrupted input data. The encoder, $\mathcal{E}(\cdot)$, creates latent representations from the input, while the decoder, $\mathcal{D}(\cdot)$, reconstructs the original data from the encoder output. The goal is to minimize the discrepancy between reconstructed and original data (Liu et al., 2022a) as follows:
(3) $\mathop{\min}_{\mathcal{D},\mathcal{E}}\ \mathcal{L}_{gen}(\mathcal{D}\circ\mathcal{E}(\omega(\mathcal{A},\mathcal{X})),\ (\mathcal{A},\mathcal{X})).$
Here, $\omega$ represents an operation like masking or perturbation. $\mathcal{D}\circ\mathcal{E}$ denotes the process of encoding and decoding to reconstruct output. Recent studies have introduced a decoder-only architecture that effectively reconstructs data without an encoder-decoder setup. This approach uses a single model (e.g., Transformer (Vaswani et al., 2017)) for reconstruction, and is commonly applied in sequential recommendation with generative learning (Sun et al., 2019). The loss function $\mathcal{L}_{gen}$ format depends on the data type, with mean square loss for continuous data and cross-entropy loss for categorical data.
2.2.3. Adversarial Learning
Adversarial learning is a training method used to generate high-quality outputs, using a generator $\mathcal{G}(\cdot)$. What sets adversarial learning apart from generative learning is the inclusion of a discriminator $\Omega(\cdot)$, which determines whether a given sample is real or generated (Gui et al., 2021; Jabbar et al., 2021). In adversarial learning, the generator aims to enhance the quality of its generated outputs in order to deceive the discriminator. Consequently, the learning objective of adversarial learning can be defined as follows:
(4) $\underset{\mathcal{G}}{\min}\ \underset{\Omega}{\max}\ \{\mathbb{E}_{x\sim P(\mathcal{A},\mathcal{X})}[\log\ {\Omega}(x)]+\mathbb{E}_{\hat{x}\sim P(\mathcal{G}(\mathcal{A},\mathcal{X}))}[\log\ (1-{\Omega}(\hat{x}))]\}.$
Here, the variable $x$ represents a real sample obtained from the underlying data distribution, while $\hat{x}$ represents a synthetic sample generated by the generator $G(\cdot)$. During the training process, both the generator and discriminator enhance their capabilities through a competitive interplay. Ultimately, the generator strives to generate high-quality outputs that can be advantageous for downstream tasks. This training approach is commonly employed in various domains, including sequential recommendation, to enhance performance.
3. Taxonomy
Figure 1. The proposed taxonomy of self-supervised learning in recommender systems.
In this section, we present our comprehensive taxonomy of self-supervised learning in recommender systems. As previously mentioned, the self-supervised learning paradigm can be categorized into contrastive learning, generative learning, and adversarial learning. Therefore, our taxonomy is built upon these three categories, providing more detailed insights into each. Our overall taxonomy is shown in Figure 1.
3.1. Contrastive Learning in Recommendation
The fundamental principle of contrastive learning (CL) involves maximizing the agreement between different views. Hence, we propose a view-centric taxonomy, which contains three key components to consider when applying contrastive learning: creating the views, pairing the views to maximize agreement, and optimizing the agreement.
3.1.1. View Creation
The created view emphasizes various data aspects for the model. It can incorporate global collaborative information to improve the recommender’s handling of global relationships (Xia et al., 2022b) or introduce random noise to enhance model robustness (Yu et al., 2022). We regard augmentation on input data (e.g., graphs, sequences, input features) (Yu et al., 2023b; Jing et al., 2023) as data-perspective view creation, and hidden feature augmentation during inference as feature-level view creation. Additionally, the model-based contrastive paradigm (Yu et al., 2023b) serves as model-level view generation. Thus, we propose a hierarchical taxonomy encompassing view-creation techniques from basic data-level to neural model-level.
  • Data-based view creation: In the realm of contrastive learning-based recommenders, diverse views are created by augmenting input data. These augmented data points are subsequently processed through deep neural recommenders. The resulting output embeddings from different views are then paired and utilized for contrastive learning. The augmentation methods for the original data vary depending on the recommendation scenario. For instance, graph-based data may employ techniques such as node/edge dropout or the addition of noisy edges (Wu et al., 2021b; Zhou et al., 2023a), while item sequences may make use of masking, cropping, and replacing (Sun et al., 2019; Zhou et al., 2020).
  • Feature-based view creation: In addition to generating views directly from the data, some methods consider conducting augmentation based on the hidden feature encoded during the models’ forward process. These representations can include node embedding during graph-based message passing or token vectors in the Transformer, for example. By applying various augmentation techniques or introducing random perturbations to the representations multiple times (Wu et al., 2021b; Yu et al., 2023a), the model’s final output can be considered as different views. One commonly used practice is to add random noise to the representations (Xie et al., 2021).
  • Model-based view creation: Both data-based and feature-based augmentation are non-adaptive since they are non-parametric. Consequently, various sophisticated methods have emerged to generate different views using neural modules. These views contain specific information based on the model design. For example, intent disentanglement neural modules can capture user intents (Ren et al., 2023b; Wu et al., 2023c), while hyperparagraph graph modules can capture global relationships (Xia et al., 2022b). In contrast, model-based view creation involves learnable parameters within the view generator that are adaptive to the learning objective and optimized during the learning process.
Figure 2. Taxonomy of Contrastive Learning in Recommender Systems.
3.1.2. Pair Sampling
The view creation process generates at least two distinct views for each sample in the data using appropriate view creation methods. The crux of contrastive learning lies in maximally aligning certain views (i.e., pulling them close) while pushing others apart. To achieve this, it’s crucial to determine the positive pair of views that should be pulled close and identify other views that form negative pairs to be pushed away. This strategy is known as pair sampling, and it primarily consists of two main pair-sampling approaches in CL-based recommendation:
  • Natural Sampling: One common approach to pair sampling is straightforward and non-heuristic, which we refer to as natural sampling. Positive pairs are formed by different views generated from the same data sample, while negative pairs are formed by views from different data samples. In cases where a central view exists for all samples, such as a global view derived from the entire graph, the local-global relationship also forms positive pairs. This approach is widely applied in most contrastive learning recommenders.
  • Score-based Sampling: Another approach to pair sampling is score-based sampling. In this approach, a module calculates scores of pairs to determine positive or negative pairs. For instance, one relevant score is the distance between two views, which can be used to form pairs (Xia et al., 2022a; Qiu et al., 2022). Alternatively, clustering can be applied on views, where positive pairs are those within the same cluster, and negative pairs are those in different clusters (Nie et al., 2022).
For an anchor view, once the positive pairs are determined, the remaining views can naturally be considered as negative views, which can be paired with the given view to create negative pairs, allowing for pushing away. Therefore, in the subsequent discussion on different methods of pair sampling, we primarily focus on the construction of positive pairs.
3.1.3. Contrastive Objective
The learning objective in contrastive learning is to maximize the mutual information between positive views, which, in turn, leads to improved performance in learning recommender models. Since directly calculating mutual information is not feasible, a tractable lower bound is commonly used as the learning objective in contrastive learning (Belghazi et al., 2018). However, there are also explicit objectives that directly pull positive views closer together.
  • InfoNCE-based Objective: InfoNCE (Oord et al., 2018), a variant of Noise Contrastive Estimation (Gutmann and Hyvärinen, 2010), has gained wide adoption as a learning objective in the field of contrastive learning for recommendation systems. The mathematical formulation of InfoNCE can be expressed as follows:
    (5) $\mathcal{L}=\mathbb{E}[-\log\frac{exp(f(\omega^{\prime}_{i},\omega^{\prime\prime}_{i}))}{\sum_{\forall i,j}exp(f(\omega^{\prime}_{i},\omega^{\prime\prime}_{j}))}]$
    Here, $f(\cdot)$ represents a critic function that calculates a score indicating the similarity between two views. The term $f(\omega^{\prime}_{i},\omega^{\prime\prime}_{i})$ corresponds to the score of positive pairs, while the term $\sum\exp(f(\omega^{\prime}_{i},\omega^{\prime\prime}_{j}))$ encompasses both the numerator and the scores of all negative pairs. By optimizing this, the $f(\cdot)$ will learn to assign higher values to positive pairs. This process aims to bring the positive pairs closer together and push the negative pairs apart.
  • JS-based Objective: In addition to using InfoNCE estimation for mutual information, the lower bound can also be estimated using the Jensen-Shannon (JS) divergence (Hjelm et al., 2018; Jing et al., 2023). The derived learning objective is akin to combining InfoNCE with a standard binary cross-entropy loss (Xia et al., 2021b), applied to positive pairs and negative pairs:
    (6) $\mathcal{L}=\mathbb{E}[-\log\sigma(f(\omega^{\prime}_{i},\omega^{\prime\prime}_{i}))]-\mathbb{E}[\log(1-\sigma(f(\omega^{\prime}_{i},\omega^{\prime\prime}_{j})))]$
    Here, $\sigma$ represents the sigmoid function used to normalize the output of the critic function. The main idea behind this optimization is to assign the label $1$ to positive pairs and 00 to negative pairs, thereby increasing the predicted value for positive pairs and enhancing the similarity between them.
  • Explicit Objective: Both InfoNCE-based and JS-based objectives maximize the estimated lower bound of mutual information to maximize mutual information itself, which is theoretically guaranteed. Additionally, there are explicit objectives, such as minimizing Mean Square Error or maximizing cosine similarity between samples within a positive pair, that directly align positive pairs. These objectives are referred to as explicit objectives.
3.2. Generative Learning in Recommendation
In generative self-supervised learning (GL), the primary objective is to maximize the likelihood estimation of the real data distribution. This allows the learned meaningful representations to capture the underlying structure and patterns in the data, which can then be utilized for downstream tasks. In our taxonomy, we consider two aspects to differentiate various recommendation methods with GL: generative learning paradigm and generation target.
Figure 3. Taxonomy of Generative Learning in Recommender Systems.
3.2.1. Learning Paradigm
In the context of recommendation, self-supervised methods employing generative learning can be classified into three paradigms: Masked Autoencoding, Variational Autoencoding and Denoised Diffusion.
  • Masked Autoencoding: In masked autoencoders, the learning process follows a mask-reconstruction approach, where the model reconstructs complete data from partial observations (He et al., 2022). In recommender systems, data can be in item sequences or user-item interaction graphs, each requiring specific masking techniques. For item sequences, certain items are randomly masked and fed into a transformer model to encode representations and reconstruct masked item features (Sun et al., 2019). For graph data, masking occurs at the edge or node level. In GraphMAE (Hou et al., 2022a), node features are masked for reconstruction, while in MaskGAE (Li et al., 2022b), edges are masked. This masking practice is common in generative self-supervised learning recommender models.
  • Variational Autoencoding: The variational autoencoder (Kingma and Welling, 2013a) is another generative approach that maximizes the likelihood estimation with theoretical guarantees. Typically, it involves mapping the input data to latent factors that follow a normal Gaussian distribution. Subsequently, the model reconstructs the input data based on sampled latent factors. Naturally, during the optimization process of VAE-based models, the learned latent factors are regularized using KL-divergence to enforce them to follow a Gaussian distribution. In self-supervised learning methods employing variational autoencoders, the latent factors may further be regulated through additional approaches, such as enhancing distinguishability through adversarial learning or contrastive learning (Wang et al., 2022e).
  • Denoised Diffusion: Denoising Diffusion is a generative model that generates new data samples by reversing a noising process. In the forward process, Gaussian noise is added to the original data in multiple steps, creating a sequence of noisy versions. In the reverse process, the model learns to remove the noise from the noisy versions, recovering the original data step by step. The model is trained to denoise the data at each step, learning to capture minor changes in the complex generation process. The diffusion model has recently been introduced into the recommendation for processing data like graphs (Jiang et al., 2024) or user interactions (Wang et al., 2023d).
3.2.2. Generation Target
What pattern of the data will be considered as the label for generation in generative learning is yet another thing that needs to be considered to bring meaningful auxiliary self-supervised signals. In general, the generation target can vary for different methods and also in different recommendation scenarios. For instance, in sequential recommendation, the generation target can be the item in the sequence (Sun et al., 2019; Sachdeva et al., 2019), in order to model the relationships within the item sequence. For recommendation with interaction graph, the generation target can be the node/edge in the graph (Xia et al., 2023; Li et al., 2023e), in order to capture high-level topological relevance in the graph.
3.3. Adversarial Learning in Recommendation
In the context of adversarial learning (AL) in recommender systems, the discriminator plays a crucial role in distinguishing between generated fake samples and real samples. Similar to generative learning, we propose our taxonomy that encompasses AL in recommender systems from both the learning paradigm and discrimination target perspectives:
Figure 4. Taxonomy of Adversarial Learning in Recommender Systems.
3.3.1. Learning Paradigm
In recommender systems, AL consists of two different paradigms, depending on whether the discriminative loss from the discriminator can be backpropagated to the generator in a differentiable way.
  • Differentiable AL The first approach involves objects represented as features in continuous space, where the gradient from the discriminator can naturally backpropagate to the generator for optimization. This method is known as differentiable adversarial learning. Here, the optimization loss aligns directly with the general learning objective in Equation (4) using a neural network-based discriminator. The discriminated objects vary depending on the method, such as latent representations from a variational autoencoder’s forward process, output features from a Transformer model, or generated user-item interaction matrices. As the quality of the generated content improves, the recommender system’s performance is simultaneously enhanced.
  • Non-differentiable AL Another approach involves discriminating the output of the recommender system, specifically the recommended items. However, since the recommended results are discrete, back-propagation becomes challenging, creating a non-differentiable situation where the gradient of the discriminator cannot be directly propagated to the generator (Wang et al., 2017; Zhao et al., 2021a). To address this, reinforcement learning with policy gradient is incorporated (Sutton et al., 1999). In this scenario, the generator acts as an agent that interacts with the environment by predicting the next item based on previously interacted items. The discriminator serves as the reward function, providing a reward signal to guide the generator’s learning. The rewards from the discriminator are defined to emphasize different factors influencing recommendation quality, and it is optimized to assign higher rewards to true samples over generated ones, guiding the generator to produce high-quality recommendations.
3.3.2. Discrimination Target
Different recommendation algorithms lead to the generation of different inputs by the generator, which are then fed to the discriminator for discrimination. This process aims to enhance the generator’s ability to generate high-quality content, thereby approaching the ground truth. The specific discrimination objectives are designed based on the particular recommendation task. For instance, in the sequential recommendation, the discriminator may be employed to discriminate generated next-items (Lv et al., 2021) or the entire sequence. In other domains, the discriminator may supervise the quality of generated data or optimize the hidden features of the model (Su et al., 2022; Zhang et al., 2021b).
4. Self-Supervised Learning for Recommendation
The study of recommendation algorithms encompasses a variety of scenarios. In this paper, we conducted a survey specifically addressing ten scenarios utilizing self-supervised learning: General Collaborative Filtering, Sequential Recommendation, Social Recommendation, Knowledge-aware Recommendation, Cross-domain Recommendation, Group Recommendation, Bundle Recommendation, Multi-behavior Recommendation, and Multi-modal Recommendation.
4.1. General Collaborative Filtering
Table 1. A summary of SSL recommendation methods in General Collaborative Filtering (Part 1). For alphabets in ”Category”, CL means Contrastive Learning; GL means Generative Learning; AL means Adversarial Learning
General Collaborative Filtering (CF)
Category Method Information Self-supervised Learning Paradigm
CL Method Year Venue View Creation Pair Sampling Contrastive Objective
Liu et al. (Liu et al., 2021c) 2021 arXiv Data-based Natural InfoNCE-based
SimpleX (Mao et al., 2021) 2021 CIKM Model-based Natural Explicit-based
EGLN (Yang et al., 2021b) 2021 SIGIR Data & Model-based Natural JS-based
SGL (Wu et al., 2021b) 2021 SIGIR Data-based Natural InfoNCE-based
BiGI (Cao et al., 2021) 2021 WSDM Data & Model-based Natural JS-based
GDCL (Zhang et al., 2022b) 2022 DASFAA Model-based Natural InfoNCE-based
DirectAU (Wang et al., 2022d) 2022 KDD Model-based Natural Explicit
SHT (Xia et al., 2022a) 2022 KDD Model-based Score-based Explicit
HCCF (Xia et al., 2022b) 2022 SIGIR Model-based Natural InfoNCE-based
RGCF (Tian et al., 2022b) 2022 SIGIR Data-based Natural InfoNCE-based
SimGCL (Yu et al., 2022) 2022 SIGIR Feature-based Natural InfoNCE-based
NCL (Lin et al., 2022b) 2022 WWW Model-based Natural InfoNCE-based
LightGCL (Cai et al., 2023) 2023 ICLR Data-based Natural InfoNCE-based
AdaGCL (Jiang et al., 2023) 2023 KDD Data-based Natural InfoNCE-based
AdvInfoNCE (Zhang et al., 2024b) 2023 NeurIPS Model-based Natural InfoNCE-based
AdaMCL (Zhu et al., 2023) 2023 SIGIR Data-based Natural InfoNCE-based
CGCL (He et al., 2023) 2023 SIGIR Model-based Natural InfoNCE-based
DCCF (Ren et al., 2023b) 2023 SIGIR Model-based Natural InfoNCE-based
uCTRL (Lee et al., 2023) 2023 SIGIR Model-based Natural Explicit
VGCL (Yang et al., 2023c) 2023 SIGIR Data-based Natural & Score-based InfoNCE-based
XSimGCL (Yu et al., 2023a) 2023 TKDE Feature-based Natural InfoNCE-based
RocSE (Ye et al., 2023a) 2023 TOIS Data-based Natural InfoNCE-based
SelfCF (Zhou et al., 2023a) 2023 TORS Data & Feature-based Natural Explicit
DENS (Lai et al., 2023) 2023 WSDM Model-based Natural Explicit
SGCCL (Li et al., 2023a) 2023 WSDM Data-based Natural InfoNCE-based
RecDCL (Zhang et al., 2024a) 2024 WWW Data & Model-based Natural Explicit
Continued on next page
Table 2. A summary of SSL recommendation methods in General Collaborative Filtering (Part 2).
General Collaborative Filtering (CF)
Category Method Information Self-supervised Learning Paradigm
GL Method Venue Year Generative Learning Paradigm Generation Target
CVAE (Li and She, 2017) 2017 KDD Variational Autoencoding User Rating & Item Content
Mult-VAE (Liang et al., 2018) 2018 WWW Variational Autoencoding User Rating to Items
MacridVAE (Ma et al., 2019) 2019 NeurIPS Variational Autoencoding User Rating to Items
RecVAE (Shenbin et al., 2020) 2020 WSDM Variational Autoencoding User Rating to Items
BiVAE (Truong et al., 2021) 2021 WSDM Variational Autoencoding User-Item Interactions
FastVAE (Chen et al., 2022a) 2022 WWW Variational Autoencoding User Rating to Items
MD-CVAE (Zhu and Chen, 2022) 2022 WWW Variational Autoencoding User Rating & Item Content
SE-VAE (Cho and Oh, 2022) 2022 WWW Variational Autoencoding User Rating to Items
CaD-VAE (Wang et al., 2023a) 2023 SIGIR Variational Autoencoding User Rating to Items
DiffRec (Wang et al., 2023d) 2023 SIGIR Denoising Diffusion User Rating to Items
GFormer (Li et al., 2023e) 2023 SIGIR Masked Autoencoding Edge in User-Item Graph
AutoCF (Xia et al., 2023) 2023 WWW Masked Autoencoding Edge in User-Item Graph
AL Method Year Venue Adversarial Learning Paradigm Discrimination Target
IRGAN (Wang et al., 2017) 2017 SIGIR Non-differentiable Item Index
CFGAN (Chae et al., 2018) 2018 CIKM Differentiable User Purchase Vector
Krishnan et al. (Krishnan et al., 2018) 2018 CIKM Non-differentiable Popular-niche Item Pair
ABinCF (Wang et al., 2019b) 2019 AAAI Differentiable Item Index
AugCF (Wang et al., 2019c) 2019 KDD Differentiable User-Item-Rating Tuple
RAGAN (Chae et al., 2019) 2019 WWW Differentiable User Purchase Vector
Conv-GCF (Yuan et al., 2020b) 2020 CIKM Differentiable Latent Interaction Map
4.1.1. Task Formulation
In General Collaborative Filtering (CF), there is no additional observed data $\mathcal{X}$, in which the model solely relies on user-item interactions $\mathcal{A}$ to generate personalized recommendations for uninteracted items.
4.1.2. Contrastive Learning in CF
Contrastive learning, as highlighted in Table 1, encompasses various SSL-based methods in collaborative filtering. These methods intelligently leverage user-item interaction data to generate diverse contrast views, facilitating effective learning in the recommendation system.
  • View Creation in CL-based collaborative filtering covers diverse paradigms. (i) Data-based View Creation. Existing works employ different augmentation techniques to create views of the data for contrastive learning. For example, Liu et al.(Liu et al., 2021c) and methods like SGL(Wu et al., 2021b) and SelfCF (Zhou et al., 2023a) use edge/node dropout to augment the user-item graph. RGCF (Tian et al., 2022b) incorporates graph structure learning to generate two different graph views, such as a denoised graph and a diversity graph. Similar ideas are also utilized in EGLN (Yang et al., 2021b). AdaMCL (Zhu et al., 2023) and SGCCL (Li et al., 2023a) create user-user and item-item graphs based on the data and then encode views for users and items. LightGCL (Cai et al., 2023) created another user-item graph based on SVD decomposition that captures global relationship modeling. Recently, RecDCL (Zhang et al., 2024a) interpolates historical and current embeddings to obtain augmented views for batch contrastive learning (ii) Feature-based View Creation. In collaborative filtering, various feature-based augmentations involve adding noise to node features during model inference. Methods such as SimGCL (Yu et al., 2022) and XSimGCL (Yu et al., 2023a) introduce controllable random noise to node embeddings during message passing, thereby creating noised views of nodes. Another approach, RocSE (Ye et al., 2023a), employs embedding perturbation techniques inspired by adversarial attacks to generate different embedding views. (iii) Model-based View Creation. Different methods also employ various neural modules to create views. On one hand, SimpleX (Mao et al., 2021), DirectAU (Wang et al., 2022d), RecDCL-FCL (Zhang et al., 2024a), CGCL (He et al., 2023), AdvInfoNCE (Zhang et al., 2024c), uCTRL (Lee et al., 2023) and XSimGCL (Yu et al., 2023a) directly utilize the encoded user/item embeddings from models as views. On the other hand, some methods incorporate carefully designed neural modules to encode views from different perspectives. For example, EGLN (Yang et al., 2021b) and BiGI (Cao et al., 2021) focus on global representation fusion. GDCL (Zhang et al., 2022b) approximates diffusion models on graphs. HCCF (Xia et al., 2022b) and SHT (Xia et al., 2022a) employ hypergraph global encoding. DCCF (Ren et al., 2023b) explores intent disentanglement. NCL (Lin et al., 2022b) and CGCL (He et al., 2023) incorporate diverse neighbor view discovery. VGCL (Yang et al., 2023c) utilizes variational latent space sampling. DENS (Lai et al., 2023) incorporates item factor encoding. Recently, AdaGCL (Jiang et al., 2023) designed a graph generative model and a graph denoising model to create adaptive contrastive views.
  • Pair Sampling in this scenario includes both natural sampling and also score-based sampling. (i) Natural Sampling. Most of the works in collaborative filtering follow natural sampling of positive samples for the anchor views to create positive pairs. There are three common situations for natural sampling in CF. Firstly, when the model encodes multiple views for one user/item, any two of these views can form a pair (Wu et al., 2021b; Cai et al., 2023; Yu et al., 2022; Xia et al., 2022b; Ren et al., 2023b). Secondly, the views from the interacted user and item can form a pair (Wang et al., 2022d; Lee et al., 2023; Zhang et al., 2024c). Lastly, by considering the relevant/irrelevant factor relationship with user-item interaction, pairs of views can be naturally formed (Lai et al., 2023). (ii) Score-based Sampling. SHT calculates edge solidity scores and samples pairs of edges for learning. VGCL clusters users/items based on embedding similarity and pairs users from the same cluster.
  • Contrastive Objective utilizes InfoNCE-based or JS-based objectives to optimize mutual information. Notably, AdvInfoNCE improves the weighting of negative pairs in InfoNCE by incorporating hardness scores learned through a min-max game. Some methods explicitly design loss functions to enhance the similarity between positive pairs. This can be achieved through cosine similarity optimization (Mao et al., 2021; Zhou et al., 2023a; Lai et al., 2023), alignment/uniformity regularization (Wang et al., 2022d; Lee et al., 2023; Zhang et al., 2024a), or by explicitly regulating the score difference between pairs (Xia et al., 2022a).
4.1.3. Generative Learning in CF
In collaborative filtering, certain self-supervised learning (SSL) methods leverage generative learning techniques (as indicated in Table 2) to reconstruct user-item interactions. This approach allows them to obtain self-supervised signals that aid in the learning process of the model.
  • Generative Learning Paradigm in self-supervised learning CF has evolved from variational autoencoding to masked autoencoding and denoised diffusion. (i) Variational Autoencoding. Different VAE-based methods in self-supervised learning CF adopt various perspectives to improve the model. For instance, CVAE (Li and She, 2017) and MD-CVAE (Zhu and Chen, 2022) incorporate item content information to regulate the collaborative latent variable modeling. RecVAE (Shenbin et al., 2020) follows the general paradigm in Mult-VAE (Liang et al., 2018) but introduces a novel composite prior distribution and a new approach to setting hyperparameters. MacridVAE (Ma et al., 2019) and CaD-VAE (Wang et al., 2023a) focus on modeling the latent factor behind interactions to capture the latent variable distribution. BiVAE (Truong et al., 2021) extends the VAE paradigm to symmetrically generate user-item interactions. FastVAE (Chen et al., 2022a) enhances model efficiency with inverted multi-index. SE-VAE (Cho and Oh, 2022) extends latent variable modeling with multi-experts and stochastic expert selection. (ii) Masked Autoencoding. Both AutoCF (Xia et al., 2023) and GFormer (Li et al., 2023e) adopt the mask-reconstruction paradigm, where they automatically identify valuable edges (user-item interactions) in the graph and reconstruct them using the learned user/item representations. (iii) Denoised Diffusion. DiffRec (Wang et al., 2023d) has introduced the diffusion model into collaborative filtering. This method leverages diffusion to denoise perturbed user ratings or latent variables, which helps mitigate the high resource costs associated with large-scale item prediction.
  • Generation Target. Both VAE-based methods (Liang et al., 2018; Ma et al., 2019; Shenbin et al., 2020; Chen et al., 2022a; Cho and Oh, 2022; Wang et al., 2023a) and diffusion-based methods (Wang et al., 2023d) share a similar auto-encoding architecture. In these methods, the user ratings to items, typically represented as vectors, serve as the generation target for the model. However, some VAE-based methods like CVAE and MD-CAE also generate item content to achieve hybrid generation. BiVAE adopts a symmetric approach by modeling the user and item latent variables, with user-item interactions as the generation target. On the other hand, in graph-based methods (Xia et al., 2023; Li et al., 2023e), valuable edges in the graph are masked and then reconstructed.
4.1.4. Adversarial Learning in CF
Collaborative filtering methods also leverage adversarial learning (as shown in Table 2) to improve recommendation indices, user ratings, or even perform data augmentation. These methods optimize the model with discriminator and enhance the overall performance of collaborative filtering systems.
  • Adversarial Learning Paradigm in collaborative filtering encompasses both non-differentiable and differentiable training approaches. IRGAN (Wang et al., 2017) introduced the idea of GANs into recommendation systems, sampling item indices from the generator for discrimination. However, the discrete nature of these item indices prevents the gradient from the discriminator from propagating to the generator for optimization. To overcome this issue, IRGAN employs policy gradient techniques, which are also adopted by Krishnan et al. (Krishnan et al., 2018). In the context of differentiable adversarial training, several methods such as CFGAN (Chae et al., 2018), RAGAN (Chae et al., 2019), and Conv-GCF (Yuan et al., 2020b) feed the generated continuous real-valued variables into the discriminator. On the other hand, ABinCF (Wang et al., 2019b) and AugCF (Wang et al., 2019c) use the Gumbel trick to approximate sampling, enabling end-to-end adversarial training.
  • Discrimination Target varies across different methods in collaborative filtering. IRGAN samples the indices of items relevant to a given user for discrimination and ABinCF follows the similar paradigm but makes the parameters of generator binary codes for efficiency. Additionally, CFGAN improves IRGAN by enabling gradient back-propagation through discriminating the user purchase vector generated by the generator. RAGAN extends this further by incorporating low ratings for items in the user purchase vector to address biased predictions. Krishnan et al. (Krishnan et al., 2018) leverage popular-niche item pairs for discrimination to improve long-tail recommendation. AugCF treats each user-item-rating tuple as a sample for the discriminator, while Conv-GCF encodes the latent interaction map of a user-item pair using a convolutional neural network for further discrimination. These methods employ different approaches to enhance recommendation performance in collaborative filtering.
4.2. Sequential Recommendation
Table 3. A summary of SSL recommendation methods in Sequential Recommendation (Part 1).
Sequential Recommendation (SeqRec)
Category Method Information Self-supervised Learning Paradigm
CL Method Year Venue View Creation Pair Sampling Contrastive Objective
S3-Rec (Zhou et al., 2020) 2020 CIKM Data-based Natural InfoNCE-based
Ma et al. (Ma et al., 2020) 2020 KDD Data & Model-based Natural InfoNCE-based
DHCN (Xia et al., 2021b) 2021 AAAI Model-based Natural JS-based
CoSeRec (Liu et al., 2021a) 2021 ArXiv Data-based Natural InfoNCE-based
CCL (Bian et al., 2021) 2021 CIKM Data-based Natural InfoNCE-based
COTREC (Xia et al., 2021a) 2021 CIKM Model-based Score-based InfoNCE-based
H2SeqRec (Li et al., 2021) 2021 CIKM Data-based Natural & Score-based InfoNCE-based
CLUE (Cheng et al., 2021) 2021 ICDM Data & Model-based Natural Explicit
MMInfoRec (Qiu et al., 2021a) 2021 ICDM Model-based Natural InfoNCE-based
SSI (Yuan et al., 2021) 2021 IJCAI Data-based Natural InfoNCE-based
ACVAE (Xie et al., 2021) 2021 WWW Feature-based Natural JS-based
CBiT (Du et al., 2022a) 2022 CIKM Data-based Natural InfoNCE-based
ContrastVAE (Wang et al., 2022e) 2022 CIKM Model-based Natural InfoNCE-based
EC4SRec (Wang et al., 2022a) 2022 CIKM Data-based Natural InfoNCE-based
MCLSR (Wang et al., 2022b) 2022 CIKM Model-based Natural InfoNCE-based
MIC (Nie et al., 2022) 2022 CIKM Data & Feature-based Score-based InfoNCE-based
TCPSRec (Tian et al., 2022a) 2022 CIKM Data-based Natural InfoNCE-based
CL4SRec (Xie et al., 2022b) 2022 ICDE Data-based Natural InfoNCE-based
DCAN-PSSL (Wang et al., 2022c) 2022 ICDE Model-based Natural Explicit
MISS (Guo et al., 2022) 2022 ICDE Model-based Score-based InfoNCE-based
GCL4SR (Zhang et al., 2022a) 2022 IJCAI Data & Model-based Natural InfoNCE-based
DCN (Lin et al., 2022a) 2022 SIGIR Data-based Natural Explicit
DuoRec (Qiu et al., 2022) 2022 WSDM Model-based Score-based InfoNCE-based
ICLRec (Chen et al., 2022b) 2022 WWW Data & Model-based Natural InfoNCE-based
TiCoSeRec (Dang et al., 2023) 2023 AAAI Data-based Natural InfoNCE-based
ECGAN-Rec (Ni et al., 2023) 2023 IPM Data & Model-based Natural Explicit
FDSA_CL (Hao et al., 2023a) 2023 TKDE Model-based Natural InfoNCE-based
ContraRec (Wang et al., 2023b) 2023 TOIS Data-based Natural InfoNCE-based
MrTransformer (Ma et al., 2023) 2023 TOIS Model-based Score-based Explicit
IOCRec (Li et al., 2023c) 2023 WSDM Data & Model-based Natural InfoNCE-based
DCRec (Yang et al., 2023b) 2023 WWW Model-based Natural InfoNCE-based
Meta-SGCL (Hao et al., 2023b) 2024 ICDE Model-based Natural InfoNCE-based
Liu et al. (Liu et al., 2024b) 2024 SIGIR Model-based Natural Explicit
ICSRec (Qin et al., 2024) 2024 WSDM Model-based Natural InfoNCE-based
Continued on next page
Table 4. A summary of SSL recommendation methods in Sequential Recommendation (Part 2).
Sequential Recommendation (SeqRec)
Category Method Information Self-supervised Learning Paradigm
GL Method Venue Year Generative Learning Paradigm Generation Target
BERT4Rec (Sun et al., 2019) 2019 CIKM Masked Autoencoding Item in Sequence
SVAE (Sachdeva et al., 2019) 2019 CIKM Variational Autoencoding Item in Sequence
PTUM (Wu et al., 2020b) 2020 EMNLP Masked Autoencoding Behavior in Sequence
PeterRec (Yuan et al., 2020a) 2020 SIGIR Masked Autoencoding Item in Sequence
U-BERT (Qiu et al., 2021b) 2021 AAAI Masked Autoencoding Words in Review
BiCAT (Jiang et al., 2021) 2021 arXiv Masked Autoencoding Item in Sequence
ShopperBERT (Shin et al., 2021) 2021 arXiV Masked Autoencoding Item in Sequence
UPRec (Xiao et al., 2021) 2021 arXiv Masked Autoencoding Item in Sequence
VSAN (Zhao et al., 2021b) 2021 ICDE Variational Autoencoding Item in Sequence
ASReP (Liu et al., 2021b) 2021 SIGIR Masked Autoencoding Item in Sequence
CBiT (Du et al., 2022a) 2022 CIKM Masked Autoencoding Item in Sequence
ContrastVAE (Wang et al., 2022e) 2022 CIKM Variational Autoencoding Item in Sequence
DiffuASR (Liu et al., 2023b) 2023 CIKM Denoised Diffusion Augmented Item Sequence
Diff4Rec (Wu et al., 2023b) 2023 MM Denoised Diffusion User-Item Interaction
MAERec (Ye et al., 2023b) 2023 SIGIR Masked Autoencoding Edge in Graph
AL Method Year Venue Adversarial Learning Paradigm Discrimination Target
AOS4Rec (Zhao et al., 2021a) 2020 IJCAI Non-differentiable Item
MFGAN (Ren et al., 2020) 2020 SIGIR Non-differentiable Feature of Sequence
SAO (Manotumruksa and Yilmaz, 2020) 2020 SIGIR Differentiable Feature of Sequence
SRecGAN (Lu et al., 2021) 2021 DASFAA Differentiable Ranking Score
SSRGAN (Lv et al., 2021) 2021 DASFAA Differentiable Item
ACVAE (Xie et al., 2021) 2021 WWW Differentiable Feature of Sequence
ECGAN-Rec (Ni et al., 2023) 2023 IPM Differentiable Feature of Sequence
4.2.1. Task Formulation
In sequential recommendation (SeqRec), user-item interactions are recorded with timestamps, establishing a temporal order. Each user has a specific temporal sequence of engaged items denoted as $s_{u}=(v_{1},v_{2},...,v_{T})$, where $T$ represents the sequence length. The objective is to predict the next item ($v_{T+1}$) based on the past item sequence. In anonymous user scenarios, only item sequences are available, known as session recommendation.
4.2.2. Contrastive Learning in SeqRec
Contrastive learning sequential recommendation comprises the majority of self-supervised SeqRec methods (as indicated in Table 3). These methods leverage the sequential data to generate diverse views through techniques such as augmentation or neural modules.
  • View Creation in SeqRec includes data-based, feature-based, and model-based methods. (i) Data-based View Creation. Given that the input data is in a sequential format, numerous methods employ sequence augmentation techniques to generate diverse views of the input item sequence. In general, early works (Zhou et al., 2020; Xie et al., 2022b; Liu et al., 2021a; Shin et al., 2021; Li et al., 2021) propose non-heuristic and random augmentation methods including Crop, Mask, Reorder, Shuffle, Substitute, and Insert to obtain different input sequences as different views. TiCoSeRec (Dang et al., 2023) goes a step further by incorporating temporal information and introducing five operators (e.g., Ti-Crop and Ti-Reorder). These operators are designed to transform the input sequence into a uniform representation while taking into account the variations in time intervals. In EC4SRec (Wang et al., 2022a), importance scores are calculated for each item in a sequence, which is then utilized to guide heuristic and explanation-driven operations for generating informative views. In addition to these methods, several data-based approaches consider constructing additional data resources such as item-item transition graphs (Zhang et al., 2022a; Yang et al., 2023b), item co-interaction (Yang et al., 2023b), and user sequences for each item (Lin et al., 2022a) to obtain diverse viewpoints. (ii) Feature-based View Creation. Feature-based methods employ various augmentation on the encoded features for view creation. For instance, ACVAE (Xie et al., 2021) utilizes variational auto-encoders to encode latent features of item sequences. It then shuffles these latent features to generate different negative views of the sequence. Additionally, MIC (Nie et al., 2022) applies random dropout on the encoded features, resulting in two distinct views of user/item representation for contrastive learning. (iii) Model-based View Creation. Model-based view creation in SeqRec encompasses two distinct approaches. The first approach involves constructing specific neural modules to encode views of interest, such as intent-aware representation (Ma et al., 2020; Chen et al., 2022b; Qin et al., 2024), hypergraph representation (Xia et al., 2021b), item attributes representation (Hao et al., 2023a), variational Transformer (Hao et al., 2023b), long-short term representations (Liu et al., 2024b) and utilizing graph neural networks and to encode embeddings from auxiliary graphs (Wang et al., 2022b; Yang et al., 2023b; Lin et al., 2022a). The second approach involves applying techniques such as random dropout (Cheng et al., 2021; Qiu et al., 2021a; Wang et al., 2022e; Qiu et al., 2022) on the model parameters to generate different outputs as distinct views.
  • Pair Sampling in SeqRec has both natural sampling and score-based heuristic sampling. (i) Natural Sampling. In most cases in SeqRec, natural sampling methods (Zhou et al., 2020; Liu et al., 2021a; Bian et al., 2021; Xie et al., 2022b; Du et al., 2022a; Wang et al., 2022a; Zhang et al., 2022a) naturally considering the generated views from the same item sequences as positive views and otherwise negative views. In addition, there are other natural pairing relations utilized in SeqRec. For instance, methods like SSI (Yuan et al., 2021), TCPSRec (Tian et al., 2022a), and DCRec (Yang et al., 2023b) construct positive pairs using various combinations, such as pairing a single item or a sub-sequence with the whole sequence it belongs to. Besides, ICSRec (Qin et al., 2024) uses clustering to generate intent prototypes and aligns the intent view with related prototypes for fine-grained intent CL. (ii) Score-based Sampling. Compared to natural sampling, score-based sampling utilizes necessary calculate to determine the positive pair. COTREC (Xia et al., 2021a) combines the last-clicked item with the predicted next item to form a positive pair. H2SeqRec (Li et al., 2021), DuoRec (Qiu et al., 2022), and MrTransformer (Ma et al., 2023) employ sequence overlap as a metric to determine positive pairs. MIC (Nie et al., 2022) first applies k-means clustering to group samples, and then samples within the same group form positive pairs. Similarly, MISS (Guo et al., 2022) employs item embedding distance as a metric to sample positive pairs with varying probabilities.
  • Contrastive Objective in SeqRec typically adopts an InfoNCE-based objective to maximize the mutual information between positive pairs. However, in methods such as DHCN (Xia et al., 2021b) and ACVAE (Xie et al., 2021), a JS-based objective is utilized for optimization. Furthermore, several algorithms incorporate explicit loss functions to contrast data samples. The fundamental concept behind these methods is to encourage the proximity of samples within positive pairs. CLUE (Cheng et al., 2021) leverages cosine similarity to minimize the distance between embeddings within a positive pair. DCAN-PSSL (Wang et al., 2022c) and ECGAN-Rec (Ni et al., 2023) aim to minimize the estimated probability distribution of the next item. DCAN-PSSL uses KL/JS-divergence minimization, while ECGAN-Rec minimizes the numerical differences between the distributions. DCN (Lin et al., 2022a) and MrTransformer (Ma et al., 2023) both minimize the squared error between two embeddings within a positive pair for optimization purposes. Liu et al. (Liu et al., 2024b) explicitly align the pairwise likelihood difference scores of two edges’ long- and short-term node embeddings for denoising.
4.2.3. Generative Learning in SeqRec
Generative learning plays a significant role in sequential recommendation (as demonstrated in Table 4). These methods primarily focus on generating sequence item data or user-item graphs to provide auxiliary signals, effectively enhancing the recommendation process.
  • Generative Learning Paradigm in SeqRec predominantly follows the approach of mask-autoencoding. For instance, BERT4Rec (Sun et al., 2019) initially attempts to randomly mask an item within the item sequence. Then, it proceeds with the encoding process using a transformer and aims to reconstruct the masked item. Given that using transformers for modeling item sequence data (Kang and McAuley, 2018) in SeqRec is a prevalent choice, the natural approach of mask-autoencoding with transformers has become a primary form in subsequent works (Wu et al., 2020b; Yuan et al., 2020a; Jiang et al., 2021; Shin et al., 2021; Xiao et al., 2021; Liu et al., 2021b; Du et al., 2022a). Furthermore, MAERec (Ye et al., 2023b) adopts the concept of mask-autoencoding from graph data (Hou et al., 2022b; Li et al., 2022c) to enhance the representation learning of items. Another notable approach in generative modeling is seen in SVAE (Sachdeva et al., 2019), which leverages variational auto-encoding as a backbone. This method utilizes recurrent neural network to encode sequential records into latent variables that follow the normal distribution and subsequently generate the next-item with a re-parametrization trick (Kingma and Welling, 2013b). This line of work is further extended by methods like VSAN (Zhao et al., 2021b) and ContrastVAE (Wang et al., 2022e) by incorporating self-attention modules or contrastive learning on latent variables. Recently, denoised diffusion has emerged as a promising paradigm in sequential recommendation (SeqRec) due to its strong generation capabilities. Both DiffuASR (Liu et al., 2023b) and Diff4Rec (Wu et al., 2023b) leverage diffusion generation to create high-quality data, which in turn enhances the training of recommendation models through data augmentation.
  • Generation Target in SeqRec with generative learning exhibits a diverse range of approaches. In general, several works (Sachdeva et al., 2019; Yuan et al., 2020a; Jiang et al., 2021; Shin et al., 2021; Xiao et al., 2021; Zhao et al., 2021b; Liu et al., 2021b; Du et al., 2022a; Wang et al., 2022e) follow the initial step of BERT4Rec, which involves masking and reconstructing the items within the sequence. PTUM considers the behavior sequence of users as the target for generation, aiming to effectively model user behavior patterns for recommendation purposes. On the other hand, U-BERT focuses on generating user reviews for items to capture user behaviors and enhance recommendations, particularly in content-insufficient domains. Recently, MAERec has incorporated the concept of mask-autoencoding from the graph domain. It utilizes this approach to generate paths in the item transition graph, effectively pre-training item embeddings for improved sequential recommendation performance. For diffusion-based methods, DiffuASR generates augmented item sequences to enrich sparse data for model training, while Diff4Rec pre-trains a diffusion model by corrupting and reconstructing user-item interactions in the latent space, which are then used to produce diverse augmentations for sparse user-item interactions.
4.2.4. Adversarial Learning in SeqRec
Adversarial learning in SeqRec (as shown in Table 4) employs various strategies, leveraging encoded features, ranking scores, and historical data to train discriminators in distinguishing between real and generated item sequences, while optimizing generators to produce realistic recommendations.
  • Adversarial Learning Paradigm in SeqRec encompasses differential and non-differential methods. Differential methods, such as SAO (Manotumruksa and Yilmaz, 2020), SRecGAN (Lu et al., 2021), ACVAE (Xie et al., 2021), and ECGAN-Rec (Ni et al., 2023), utilize continuous operations for gradient propagation and avoid discrete sampling. SSRGAN (Lv et al., 2021) employs Gumble-Softmax for differentiable sampling to overcome gradient blocking. In non-differential methods, reinforcement learning optimizes the generator’s parameters after generating items through non-differentiable sampling and feeding them to the discriminator. AOS4Rec (Zhao et al., 2021a) employs the WGAN (Arjovsky et al., 2017) concept for simultaneous optimization and the actor-critic algorithm (Konda and Tsitsiklis, 1999) for stable training, while MFGAN (Ren et al., 2020) uses policy gradients to improve generator predictions using the discriminator’s score as a reward.
  • Discrimination Target exhibits a wide range of diversity within SeqRec when employing adversarial learning. Various approaches, such as MFGAN, SAO, ACVAE, and ECGAN-Rec, employ encoded features derived from the item sequence, which are then fed into the discriminator. The discriminator’s role is to determine whether the item sequence is real or generated by the generator. Besides, the discrimination target in SRecGAN is the ranking score of items, where the score difference between positive and negative items generated by the generator deceives the discriminator. The discriminator is also optimized using ground truth data with an ideal score difference, aiming to force the generator to assign higher values to positive pairs. In AOS4Rec, the discriminator calculates the score by considering the history item along with the sampled next-item. It assigns a higher value to the real item sequence. A similar concept is demonstrated in SSRGAN, but it utilizes differential sampling to circumvent the involvement of reinforcement learning.
4.3. Social Recommendation
Table 5. A summary of SSL recommendation methods in Social Recommendation and Knowledge-aware Recommendation.
Social Recommendation (SocRec)
Category Method Information Self-supervised Learning Paradigm
CL Method Year Venue View Creation Pair Sampling Contrastive Objective
KCGN (Huang et al., 2021) 2021 AAAI Model-based Natural JS-based
SMIN (Long et al., 2021) 2021 CIKM Model-based Natural JS-based
SEPT (Yu et al., 2021a) 2021 KDD Data-based Score-based InfoNCE-based
MHCN (Yu et al., 2021b) 2021 WWW Data-based Natural Explicit
DcRec (Wu et al., 2022a) 2022 CIKM Data-based Natural InfoNCE-based
SDCRec (Du et al., 2022b) 2022 SIGIR Model-based Score-based InfoNCE-based
DUAL (Tao et al., 2022a) 2022 TCSS Data-based Score-based InfoNCE-based
DSL (Wang et al., 2023c) 2023 IJCAI Model-based Natural Explicit
ReACL (Ji et al., 2023) 2023 Inf. Sci. Data-based Score-based InfoNCE-based
HGCL (Chen et al., 2023) 2023 WSDM Model-based Natural InfoNCE-based
AL Method Venue Year Adversarial Learning Paradigm Discrimination Target
Adit et al. (Krishnan et al., 2019) 2019 CIKM Non-differentiable Sampled User-User Pair
RSGAN (Yu et al., 2019) 2019 ICDM Differentiable Generated Item
DASO (Fan et al., 2019) 2019 IJCAI Non-differentiable Sampled User-User/Item Pair
ESRF (Yu et al., 2020) 2022 TKDE Differentiable Sampled Neighbor Users
Knowledge-aware Recommendation (KnoRec)
Category Method Information Self-supervised Learning Paradigm
CL Method Year Venue View Creation Pair Sampling Contrastive Objective
CKER (Pan and Chen, 2021) 2022 Mathematics Model-based Natural InfoNCE-based
KGIC (Zou et al., 2022b) 2022 CIKM Data-based Score-based InfoNCE-based
KGCL (Yang et al., 2022) 2022 SIGIR Data-based Natural InfoNCE-based
MCCLK (Zou et al., 2022a) 2022 SIGIR Model-based Natural InfoNCE-based
KRec-C2 (Peng et al., 2023) 2023 DASFAA Model-based Natural InfoNCE-based
ML-KGCL (Chen and Xie, 2023) 2023 DASFAA Model-based Natural InfoNCE-based
HiCON (Wu et al., 2023a) 2023 ICME Data-based Natural InfoNCE-based
KACL (Wang et al., 2023e) 2023 WSDM Model-based Natural InfoNCE-based
GL Method Venue Year Generative Learning Paradigm Generation Target
KGRec (Yang et al., 2023a) 2023 KDD Maksed Autoencoding Triplets in the KG
DiffKG (Jiang et al., 2024) 2024 WSDM Denoising Diffusion Triplets in the KG
4.3.1. Task Formulation
In social recommendation (SocRec), recommender systems are enhanced with side information that unveils the social relationships among users. The auxiliary observed data $\mathcal{X}$ in social recommendation is commonly represented as a user-user interaction graph $\mathcal{G}_{soc}$. In this graph, each interaction indicates the existence of a social relationship (e.g., friendship) between two users. The objective in SocRec is similar to general collaborative filtering: to recommend unexplored items to users by considering both collaborative and social information.
4.3.2. Contrastive Learning in SocRec
Contrastive learning has become a significant approach in social recommendation, as demonstrated in Table 5. By leveraging social relationship data, various contrastive views are generated, and mutual information is optimized through carefully designed pair-sampling strategies and objective functions.
  • View Creation in SocRec with CL can be categorized as data-based or model-based. The model-based approach emphasizes designing neural modules to encode features from social and collaborative data for contrastive views. Examples include KCGN (Huang et al., 2021) using behavior and temporal-aware networks, and SMIN (Long et al., 2021) and SDCRec (Du et al., 2022b) employing heterogeneous GNNs while considering social relationships. DSL (Wang et al., 2023c) utilizes GNNs for encoding distinct social-aware and collaborative-aware user embeddings. Conversely, data-based methods augment original graphs to generate diverse context-specific data. SEPT (Yu et al., 2021a) creates three graph views (preference, friend, and sharing) from user-item and user-user graphs, while MHCN (Yu et al., 2021b) generates hypergraphs among users using triangle motifs and applies hypergraph convolution. DcRec (Wu et al., 2022a) diversifies views through edge addition and node/edge dropout, and ReACL (Ji et al., 2023) constructs additional graphs using relationship-aware augmentation. In HGCL (Chen et al., 2023), a personalized knowledge transfer meta network encodes enhanced user and item representations for one view, while a lightweight GCN encodes another view for users/items.
  • Pair Sampling in SocRec is designed based on the encoded nature of the created views. In natural sampling methods, a user or item may have multiple views. For instance, KCGN and MHCN pair node embeddings with their encoded sub-graph or hypergraph embeddings, considering the local-global relationship. Furthermore, several methods (Long et al., 2021; Wu et al., 2022a; Wang et al., 2023c; Chen et al., 2023) encode multiple views for each node using graph augmentation or neural model design, treating the views of the same node as positive pairs. For score-based sampling, SEPT, SDCRec and ReACL label positive pairs by considering the representation similarly, and DUAL pre-computed a link-score for each edge in the graph based on the degree of the node, which is then utilized to choose positive pairs.
  • Contrastive Objective in these methods commonly relies on InfoNCE-based (Yu et al., 2021a; Wu et al., 2022a; Du et al., 2022b; Tao et al., 2022a; Ji et al., 2023; Chen et al., 2023) or JS-based objectives (Huang et al., 2021; Long et al., 2021). Notably, ReACL (Ji et al., 2023) introduces an Aug-InfoNCE contrastive loss, leveraging an expanded set of positive samples to improve consistency among similar nodes and enhance the framework’s generalization capability. In explicit optimization methods, MHCN employs a pair-wise ranking loss to maximize the mutual information between representations within a pair. On the other hand, DSL utilizes hinge loss to optimize the pair score, which proves beneficial in reducing the negative influence caused by noisy edges.
4.3.3. Adversarial Learning in SocRec
Adversarial learning in social recommendation, as demonstrated in Table 5, is frequently utilized to extract clean and representative social relations from the user-user graph in SocRec.
  • Adversarial Learning Paradigm. Adit et al. (Krishnan et al., 2019) proposed a modular adversarial approach to prevent model collapse in recommender systems with user-level social links. Their method utilized generative adversarial learning and a discriminator to differentiate sampled user pairs from both the interest and social domains. The optimization of parameters involved policy gradient due to the non-differentiable sampling operation. Similarly, DASO (Fan et al., 2019) leveraged the sampled user-user or user-item pairs from the generator and utilized policy gradient to optimize the parameters. RSGAN (Yu et al., 2019) uses the Gumbel-softmax technique for differentiable sampling and generates items based on social links. The discriminator in RSGAN aims to assign higher scores to both ground truth and socially aware items compared to negative items, following the social BPR optimization. In ESRF (Yu et al., 2020), a concrete selector layer is utilized to avoid non-differentiable sampling for neighbor users for discrimination.
  • Discrimination Target. In social recommendation with adversarial learning, the discrimination target aims to combine social information and collaborative information in a regularized manner. Adit et al.(Krishnan et al., 2019), RSGAN(Yu et al., 2019), and ESRF (Yu et al., 2020) all adopt a similar approach of obtaining user-user/item pairs through sampling from the generator. They then utilize the discriminator to enhance the quality of the generated samples. Notably, RSGAN (Yu et al., 2019) employs the generated items for discrimination, thereby improving the generator’s ability to generate more accurate items based on the observed ground truth items.
4.4. Knowledge-aware Recommendation
4.4.1. Task Formulation
Items in reality have diverse attributes and labels, which can form a knowledge graph (KG) when connected to corresponding items. Knowledge-aware recommendation (KnoRec) leverages this external knowledge to enhance recommendations, represented as $\mathcal{G}_{k}=(h,r,t)$, where $h$ and $t$ are knowledge entities, and $r$ is their semantic relationship (e.g., book, is written by, author). The knowledge graph $\mathcal{G}_{k}$ serves as external information to improve collaborative filtering performance, similar to social recommendation.
4.4.2. Contrastive Learning in KnoRec
Contrastive learning in knowledge-aware recommendation, as shown in Table 5 merges knowledge graphs with user-item graphs, employing various encoding, graph construction, and augmentation techniques to generate diverse perspectives for improved performance.
  • View Creation in KnoRec involves the exploration of leveraging KGs in conjunction with the existing user-item graph to construct diverse perspectives. CKER (Pan and Chen, 2021) employs light-weight graph convolution and relation-aware convolution to encode interaction-aware and knowledge-aware item representations as two distinct views. KGIC (Zou et al., 2022b) utilizes the user-item interaction graph and KG to construct local and non-local graphs, encoding different views for nodes based on these two graphs. KGCL (Yang et al., 2022) and ML-KGCL (Chen and Xie, 2023) both employ graph augmentation, such as edge dropout, twice to create contrastive views. MCCLK (Zou et al., 2022a) first builds an item-item graph as the semantic view using relation-aware GNN, and then encodes collaborative and semantic representations for each node for contrastive learning. KRec-C2 (Peng et al., 2023) incorporates a category-level aggregation representation layer to obtain category-level signal features, which are contrasted with the original category features derived from the embedding function. HiCON (Wu et al., 2023a) leverages meta-paths to construct a high-order graph that encodes representations of users and items, contrasting them with representations obtained from the original interaction graph. KACL (Wang et al., 2023e) encodes the interaction view using an augmented user-item graph and KG-aware item representations based on the KG, providing another view for contrastive learning.
  • Pair Sampling. Generally, in these methods, multiple views are created for each user and item node in the graph using various techniques (Zou et al., 2022b; Yang et al., 2022; Wang et al., 2023e). These views naturally form positive pairs for contrastive learning. In KGIC (Zou et al., 2022b), the distance between nodes on the local/non-local graph is used to select positive node pairs, enabling inter-graph and intra-graph interactive contrastive learning.
  • Contrastive Objective. In all the mentioned methods, the contrastive objective is to optimize the mutual information between different views using the InfoNCE-based lower bound.
4.4.3. Generative Learning in KnoRec
Generative learning also emerges as a novel SSL paradigm to further enhance the performance of KnoRec (Table 5). The intrinsic idea behind generative learning is to discover valuable knowledge triplets through generative tasks while mitigating the negative impact (Liu et al., 2020; Xiao et al., 2015) caused by noisy or irrelevant knowledge.
  • Generative Learning Paradigm in KnoRec encompasses both masked autoencoding and denoising diffusion approaches. In KGRec (Yang et al., 2023a), a criterion score is initially computed for each knowledge triplet to identify the most valuable rationale triplets. Subsequently, the masking-and-reconstructing paradigm is employed to reconstruct these important triplets using node embeddings in the KG. This enables the encoding of information brought by these crucial triplets into the node representation, thereby enhancing recommendation performance. DiffKG (Jiang et al., 2024) employs the concept of denoising diffusion by adding Gaussian noise to the triplets in the KG and subsequently removing it. This approach effectively eliminates the inherent noise present in the knowledge graph, providing a cleaner and augmented KG that can be utilized for inference and recommendation purposes.
  • Generation Target in both KGRec and DiffKG encompasses all the triplets present in the knowledge graph. This is because the original interactions within the KG may contain significant amounts of noise and redundant information that are irrelevant for recommendation (Yang et al., 2022). The generative learning process effectively generates valuable interactions within the KG through self-supervised learning, thereby achieving denoising.
4.5. Cross-domain Recommendation
Table 6. A summary of SSL recommendation methods in Cross-domain Recommendation.
Cross-domain Recommendation (CroRec)
Category Method Information Self-supervised Learning Paradigm
CL Method Year Venue View Creation Pair Sampling Contrastive Objective
PCRec (Wang et al., 2021) 2021 CogMI Data-based Natural InfoNCE-based
C2DSR (Cao et al., 2022a) 2022 CIKM Data-based Natural JS-based
SASS (Zhang et al., 2022c) 2022 CIKM Model-based Natural InfoNCE-based
CDRIB (Cao et al., 2022b) 2022 ICDE Model-based Natural JS-based
CLCDR (Chen et al., 2022c) 2022 ICONIP Data-based Natural InfoNCE-based
CCDR (Xie et al., 2022a) 2022 KDD Data & Model-based Natural InfoNCE-based
SITN (Sun et al., 2023) 2023 AAAI Model-based Natural InfoNCE-based
CATCL (Xiao et al., 2023) 2023 DASFAA Feature-based Natural InfoNCE-based
DCCDR (Zhang et al., 2023b) 2023 DASFAA Model-based Natural InfoNCE-based
DR-MTCDR (Guo et al., 2023a) 2023 TOIS Data-based Natural InfoNCE-based
GL Method Venue Year Generative Learning Paradigm Generation Target
RA/SA-VAE (Salah et al., 2021) 2021 RecSys Variational Autoencoding User Rating to Items
VDEA (Liu et al., 2022c) 2022 SIGIR Variational Autoencoding User Rating to Items
DiffCDR (Xuan, 2024) 2024 arXiv Denoising Diffusion User Feature
AL Method Venue Year Adversarial Learning Paradigm Discrimination Target
Su et al. (Su et al., 2022) 2022 CIKM Differentiable Domain Item Feature
RecGURU (Li et al., 2022d) 2022 WSDM Differentiable Domain User Feature
ACLCDR (Hsu et al., 2023) 2023 TKDD Non-Differentiable Augmented Interaction Matrix
DA-CDR (Zhang et al., 2021b) 2023 TKDE Differentiable Domain User/Item Feature
DA-DAN (Guo et al., 2023b) 2023 TOIS Differentiable Domain User Feature
4.5.1. Task Formulation
Cross-domain recommendation (CroRec) transfers learned user preferences from a source domain to improve recommendations in a target domain, each containing different items based on domain-specific criteria. The items are divided into source domain set $\mathcal{V}_{s}$ and target domain set $\mathcal{V}_{t}$, with interaction matrices $\mathcal{A}_{s}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{V}_{s}|}$ and $\mathcal{A}_{t}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{V}{t}|}$, where $|\mathcal{U}|$ represents user count. Recommendations for users with uninteracted items ($v\in\mathcal{V}_{t}$) are made using a function $f$ that calculates preference scores ($y_{u,v}$) based on collaborative information across domains. CroRec can also involve multiple domains and be enhanced by incorporating sequential information (Cao et al., 2022a; Sun et al., 2023).
4.5.2. Contrastive Learning in CroRec
Contrastive learning is crucial in SSL for cross-domain recommendation methods (Table 6), as it enables knowledge transfer from the target to the source domain by treating user interaction outcomes from different domains as distinct views. This section investigates diverse contrastive view creation methods using data from different domains and techniques for pair sampling and optimization in contrastive learning.
  • View Creation. For data-based methods, PCRec (Wang et al., 2021) generates two distinct views using random walks on the source domain graph and encodes representations for contrastive pre-training. C2DSR (Cao et al., 2022a) creates single-domain and cross-domain sequences, pooling their prototype embeddings to form two positive views for contrastive learning. CLCDR (Chen et al., 2022c) replaces the BPR loss with a contrastive loss, creating views and positive pairs based on interaction data. DR-MTCDR (Guo et al., 2023a) employs node/edge dropping for graph augmentation, while CCDR (Xie et al., 2022a) uses a sub-graph-based data augmentation technique for view creation in its intra-domain CL. For model-based view creation, several methods consider representations learned by neural modules as distinct views. SASS (Zhang et al., 2022c) and CDRIB (Cao et al., 2022b) leverage contrastive loss to maximize differences between views of the same node, while CCDR (Xie et al., 2022a) uses neural-encoded representations of the same nodes from different domain data for inter-domain CL. SITN (Sun et al., 2023) encodes two user-specific sequences from source and target domains as distinct views and employs a clustering technique for an additional view. DCCDR (Zhang et al., 2023b) projects original representations into domain-specific and domain-invariant representations, maximizing mutual information between domain-invariant representations of users from two domains. For feature-based methods, CATCL (Xiao et al., 2023) adopts a feature-based view creation approach, similar to SimGCL, by adding noise to the node features during inference. This noise addition process generates diverse views that can be utilized for learning purposes.
  • Pair Sampling relies on natural sampling in these works, where multiple views are created for each data object. Positive pairs are formed by two views of the same object, while views from other objects serve as negative samples. Encoded representations from multiple views for the same user/item node can be paired, originating from different domains (Zhang et al., 2022c; Cao et al., 2022b; Xie et al., 2022a) or generated through data augmentation techniques (Wang et al., 2021; Xiao et al., 2023). Additionally, these representations can be encoded by specifically designed neural modules (Wang et al., 2021; Zhang et al., 2023b).
  • Contrastive Objective within CL-based cross-domain recommendation methods typically employs the InfoNCE-based objective to optimize the mutual information between positive views (Wang et al., 2021; Zhang et al., 2022c; Chen et al., 2022c; Xie et al., 2022a; Sun et al., 2023; Xiao et al., 2023; Zhang et al., 2023b; Guo et al., 2023a). Additionally, some methods utilize the JS-based lower bound for optimization. For instance, C2DSR (Cao et al., 2022a) corrupts the prototype representation to generate negative views, while CDRIB (Cao et al., 2022b) derives the JS-based contrastive term based on the information bottleneck regularization theory.
4.5.3. Generative Learning in CroRec
In CroRec, generative learning (as shown in Table 6) employs variational autoencoding and diffusion models to align and transfer knowledge between the source and target domains. The core idea is to facilitate knowledge sharing and adaptation across domains.
  • Generative Learning Paradigm in CroRec begins by employing deep generative latent variable models (i.e., variational auto-encoders) to encode the latent space of user preferences. Subsequently, it transitions to the denoising diffusion model to facilitate knowledge transfer across domains. Salah et al. propose RA/SA-VAE (Salah et al., 2021) to simultaneously fit the target observations and align the hidden space encoded by VAE with the source latent space. They introduce both rigid alignment and soft alignment techniques with varying degrees of user preference alignment. Furthermore, VDEA (Liu et al., 2022c) extends the alignment of latent spaces using variational autoencoding at both local and global levels, allowing for the exploitation of domain-invariant features across different domains, including both overlapped and non-overlapped users. Recently, DiffCDR (Xuan, 2024) leveraged denoising probabilistic models (DPMs) to process noisy data and generate denoised results, effectively transferring data between domains.
  • Generation Target. RA/SA-VAE and VDEA both adopt the user rating reconstruction paradigm to optimize the variational model. In these models, the generation target is the user ratings for items, represented as vectors in the model. The generated ratings include ratings for uninteracted items, which are subsequently used for ranking and recommendation. On the other hand, DiffCDR focuses on generating user features in the target domain by reversing the diffusion process conditioned on the corresponding user’s information in the source domain. These generated user features in the target domain are then utilized for recommendation purposes.
4.5.4. Adversarial Learning in CroRec
Adversarial learning in Table 6 employs adversarial domain adaptation to generate domain-independent features, encoding domain-invariant user interaction preferences for effective knowledge transfer. Some methods also utilize adversarial samples to improve model learning.
  • Adversarial Learning Paradigm. Several methods employ adversarial learning for cross-domain recommendation. For instance, Su et al. (Su et al., 2022) use a discriminator to identify domain-specific features, with gradients propagating to optimize the feature generator. Similarly, RecGURU (Li et al., 2022d) feeds encoded user representations from both domains into a discriminator for classification, enabling differentiable adversarial learning. Moreover, ACLCDR (Hsu et al., 2023) utilizes Deep Q-Learning to design a generator that enhances the interaction matrix with fake interactions, optimizing it using reward signals from downstream tasks. In addition, DA-CDR (Zhang et al., 2021b) encodes domain features of users and items to fool the discriminator, ensuring effective domain knowledge transfer through adversarial learning using a gradient reversal layer (GRL) for differentiable training. Furthermore, DA-DAN (Guo et al., 2023b) incorporates adversarial domain adaptation into unsupervised non-overlapping CroRec, with the discriminator classifying the domain of encoded user features for overall differentiable training.
  • Discrimination Target. Su et al. (Su et al., 2022)’s domain adaptation method uses features from source and target domains to discriminate between them, forcing the generator to encode domain-invariant user preferences, thereby confusing the discriminator. RecGURU adopts a similar approach, discriminating encoded user representations from both domains to generalize user representation learning across domains. In ACLCDR, the augmented interaction matrix from DQN is used for inference in subsequent models, with the recommendation model acting as an implicit discriminator. DA-CDR employs a discriminator to classify domain features of users/items, encouraging the generator to encode additional domain-shared information for cross-domain recommendation. Lastly, DA-DAN encodes user-specific item sequence representations as domain features for the discriminator’s classification task, enabling the learning of domain-invariant features for recommendation.
4.6. Bundle Recommendation
Table 7. A summary of SSL recommendation methods in Bundle Recommendation and Group Recommendation.
Bundle Recommendation (BunRec)
Category Method Information Self-supervised Learning Paradigm
CL Method Year Venue View Creation Pair Sampling Contrastive Objective
MIDGN (Zhao et al., 2022) 2022 AAAI Model-based Natural InfoNCE-based
CrossCBR (Ma et al., 2022) 2022 KDD Data & Feature-based Natural InfoNCE-based
GL Method Venue Year Generative Learning Paradigm Generation Target
DGMAE (Ren et al., 2023a) 2023 SIGIR Maksed Autoencoding Edge in User-Item Graph
Group Recommendation (GroRec)
Category Method Information Self-supervised Learning Paradigm
CL Method Year Venue View Creation Pair Sampling Contrastive Objective
GroupIM (Sankar et al., 2020) 2020 SIGIR Model-based Natural JS-based
HHGR (Zhang et al., 2021a) 2021 CIKM Data-based Natural JS-based
CubeRec (Chen et al., 2022d) 2022 SIGIR Model-based Score-based Explicit
SGGCF (Li et al., 2023d) 2023 WSDM Model-based Natural InfoNCE-based
4.6.1. Task Formulation
In Bundle Recommendation (BunRec), the auxiliary observed data $\mathcal{X}$ includes the item-bundle affiliation, where a bundle $b\in\mathcal{B}$ represents a set of items for recommendation. Therefore, we have two interaction matrices: the user-bundle interaction matrix $\mathcal{A}_{U-B}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{B}|}$ and the user-item interaction matrix $\mathcal{A}_{U-I}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{I}|}$. Additionally, we use the matrix $\mathcal{A}_{I-B}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{B}|}$ to record the item-bundle affiliation. The overall objective is to recommend uninteracted bundles to each user by predicting their preference scores $y_{u,b}$.
4.6.2. Contrastive Learning in BunRec
In bundle recommendation with contrastive learning (Table 7), the relationship between items and bundles, along with user-item interactions, is utilized to form user-bundle relationships. Various methods are then designed to create views based on these relationships.
  • View Creation. MIDGN (Zhao et al., 2022) employs neural models to encode distinct representations for each user and bundle, treating them as separate views. It incorporates a graph disentangle module to encode intent-aware representations and utilizes a user-bundle graph to encode cross-view representations with LightGCN. Additionally, CrossCBR (Ma et al., 2022) initially builds the user-bundle graph by leveraging the user-item and bundle-item graphs. Subsequently, it encodes user and bundle representations using graph augmentation and embedding augmentation techniques in the bundle and item views. As a result, each user and bundle node possesses two views for contrastive learning, similar to the approach adopted in MIDGN.
  • Pair Sampling. Both MIDGN and CrossCBR encode two distinct views for each user and bundle in the dataset. As a result, the two views of a specific user/bundle node naturally form the positive pair, while other views from different nodes serve as negative samples for contrast comparison.
  • Contrastive Objective. MIDGN and CrossCBR both utilize the InfoNCE-based objective to optimize their models through self-supervised learning signals.
4.6.3. Generative Learning in BunRec
In BunRec, generative learning (Table 7) has adopted the concept of a graph-masked auto-encoder. The approach (i.e., DGMAE (Ren et al., 2023a)) is illustrated as follows.
  • Generative Learning Paradigm. In DGMAE, a teacher GNN model is first trained using user-bundle interactions, and then its knowledge is distilled into a student graph-masked auto-encoder model. This learning paradigm in DGMAE follows the principles of mask autoencoding.
  • Generation Target. DGMAE employs an adaptive masking strategy on the prominent edges within the user-item graph for the purpose of reconstruction. Specifically, edges with lower sparsity scores are prioritized for masking in DGMAE, increasing the reconstruction challenge and improving the model’s robustness against sparsity.
4.7. Group Recommendation
4.7.1. Task Formulation
The goal of Group Recommendation (GroRec) is to recommend items to a group denoted as $o\in\mathcal{O}$, which represents a set of users. In addition to user-item interactions, there are also group-item interactions represented by $\mathcal{A}_{O-I}\in\mathbb{R}^{|\mathcal{O}|\times|\mathcal{I}|}$ and group-user affiliations represented by $\mathcal{A}_{O-U}\in\mathbb{R}^{|\mathcal{O}|\times|\mathcal{U}|}$. The recommender model will predict a score $y_{o,v}$ for each group-item pair $(o,v)$ for recommendation.
4.7.2. Contrastive Learning in GroRec
In GroRec, contrastive learning is vital in SSL methods (Table 7), utilizing diverse user-item and user-group relationships to create multiple views for each node, enabling effective training.
  • View Creation. GroupIM (Sankar et al., 2020) pioneers the use of mutual information maximization, encoding user preference and group representations as views for contrastive learning. Meanwhile, HHGR (Zhang et al., 2021a) constructs fine- and coarse-grained user-level hypergraphs, encoding augmented views of user nodes at different granularities. Moreover, CubeRec (Chen et al., 2022d) transforms the recommendation scenario using a hypercube framework, encoding user and item representations with LightGCN and treating each user representation as a view for pairing and learning. Lastly, SGGCF (Li et al., 2023d) adopts a data-based view creation approach, constructing a group-user-item graph, augmenting it with node/edge dropout, and encoding diverse views for each node.
  • Pair Sampling. In BunRec, users are linked with user groups, forming positive pairs between user and group representations. GroupIM maximizes mutual information (MI) between these representations. In HHGR and SGGCF, each graph node has multiple augmented views, which are naturally treated as positive pairs, while views of other nodes are considered negative samples. CubeRec constructs a group hypercube based on user representations, identifying the intersection between two group hypercubes. Each overlapping user is paired with the hypercube as a positive pair, and other users are considered negative samples. The distance between each user and the hypercube is then calculated for contrastive optimization.
  • Contrastive Objective. GroupIM utilizes a noise-contrastive objective with a binary cross-entropy loss, specifically a JS-based objective which maximizes the mutual information (MI) between positive pairs. On a related note, HHGR also employs a similar JS-based optimization approach. In CubeRec, an explicit margin loss is utilized to ensure that the distance-to-hypercube of positive pairs is smaller than that of negative pairs. For SGGCF, it adopts the InfoNCE-based objective to achieve the contrastive learning process.
4.8. Multi-behavior Recommendation
Table 8. A summary of SSL recommendation methods in Multi-behavior Recommendation and Multi-modal Recommendation.
Multi-behavior Recommendation (MbeRec)
Category Method Information Self-supervised Learning Paradigm
CL Method Year Venue View Creation Pair Sampling Contrastive Objective
HMG-CR (Yang et al., 2021a) 2021 ICDM Data & Model-based Natural InfoNCE-based
S-MBRec (Gu et al., 2022) 2021 ICDM Model-based Score-based InfoNCE-based
MMCLR (Wu et al., 2022b) 2022 DASFAA Model-based Natural Explicit
CML (Wei et al., 2022) 2022 WSDM Model-based Natural InfoNCE-based
IICL (Li et al., 2023b) 2023 DASFAA Model-based Natural & Score-based InfoNCE-based
MixMBR (Qiao et al., 2023) 2023 DASFAA Model-based Natural InfoNCE-based
TMCL (Xuan and Li, 2023) 2023 DASFAA Model-based Natural InfoNCE-based
RCL (Wei et al., 2023b) 2023 RecSys Model-based Natural InfoNCE-based
MBSSL (Xu et al., 2023) 2023 SIGIR Data & Model-based Natural InfoNCE-based
KMCLR (Xuan et al., 2023) 2023 WSDM Data & Model-based Natural InfoNCE-based
GL Method Venue Year Generative Learning Paradigm Generation Target
VCGAE (Liu et al., 2023a) 2023 ICDM Variational Autoencoding Target Behavior Graph
BVAE (Rao et al., 2023) 2023 RecSys Variational Autoencoding User Interaction Vector
Multi-modal Recommendation (MmoRec)
Category Method Information Self-supervised Learning Paradigm
CL Method Year Venue View Creation Pair Sampling Contrastive Objective
Liu et al. (Liu et al., 2022b) 2022 ICMR Model-based Natural JS-based
CMI (Li et al., 2022a) 2022 SIGIR Data-based Natural InfoNCE-based
MMGCL (Yi et al., 2022) 2022 SIGIR Data-based Natural InfoNCE-based
SLMRec (Tao et al., 2022b) 2022 TMM Data-based Natural InfoNCE-based
MICRO (Zhang et al., 2022d) 2023 TKDE Model-based Natural InfoNCE-based
BM3 (Zhou et al., 2023b) 2023 WWW Feature-based Natural Explicit
MMSSL (Wei et al., 2023a) 2023 WWW Model-based Natural InfoNCE-based
GL Method Venue Year Generative Learning Paradigm Generation Target
MVGAE (Yi and Chen, 2021) 2022 TMM Variational Autoencoding Edge in Graph
AL Method Venue Year Adversarial Learning Paradigm Discrimination Target
MMSSL (Wei et al., 2023a) 2023 WWW Differentiable User Interaction to Items
4.8.1. Task Formulation
In Multi-behavior Recommendation (MbeRec), the user-item interaction is extended to incorporate behavior heterogeneity, resulting in a 3D tensor denoted as $\mathcal{A}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{I}|\times|\mathcal{B}|}$, where =$\mathcal{B}=\{b_{1},...,b_{K}\}$ depicts the set of different types of behaviors (e.g., page view, add-to-cart). The MbeRec aims to provide item recommendations for the target behavior (e.g., purchase) by leveraging the diverse behavior information.
4.8.2. Contrastive Learning in MbeRec
In MbeRec, contrastive learning leverages multi-behavior user-item interactions to create diverse behavior-specific views for each data sample. These views, which differ among various methods, are ideal for contrastive learning. The methods are listed in Table 8.
  • View Creation. Various methods employ model-based view creation and data-level view augmentation. HMG-CR (Yang et al., 2021a) uses hyper meta-paths to construct multiple hyper meta-graphs, encoding diverse behavior embeddings as views for each user node using distinct graph encoders. S-MBRec (Gu et al., 2022) partitions the multi-behavior graph into sub-graphs, encoding multiple GCN embeddings for each user/item as views. MMCLR (Wu et al., 2022b) encodes sequence and graph views for each user under different behaviors, obtaining a fusion view by combining these representations. CML (Wei et al., 2022) and IICL (Li et al., 2023b) leverage a behavior-aware graph neural network to encode user embeddings for each behavior, treating each behavior type as a separate view. MixMBR (Qiao et al., 2023) encodes behavior-specific embeddings and views through behavior message passing and model-parameter mixup. TMCL (Xuan and Li, 2023) utilizes temporal information to encode diverse behavior embeddings as distinct views through temporal-aware global learning. RCL (Wei et al., 2023b) encodes behavior-specific short-term and long-term interest embeddings as separate views, combining them to create a multi-behavior pattern view. MBSSL (Xu et al., 2023) employs behavior-aware GNNs to encode views for inter-behavior contrastive learning, and utilizes edge dropout to obtain augmented views for intra-behavior contrastive learning. KMCLR (Xuan et al., 2023) encodes behavior-specific views and employs knowledge-based data augmentation to obtain KG-aware views for each item for pairing.
  • Pair Sampling. In HMG-CR, user behavior embeddings form positive pairs, with embeddings from other graphs being negative pairs. S-MBRec calculates similarity scores based on interaction nodes to determine positive user/item pairs. MMCLR, CML, MixMBR, TMCL, MBSSL and KMCLR treat different behavior views of the same node as positive pairs, and others as negative samples. In MBSSL, user similarity scores are computed to detect false negative samples and filter nagative pairs. Moreover, KMCLR also pairs two knowledge-augmented views of the same item for knowledge-aware contrastive learning. IICL uses intra- and inter-behavior contrastive learning, where even-layered embeddings are positive pairs in intra-behavior CL, and different behaviors for the same user are positive pairs in inter-behavior CL. Lastly, in RCL, the fused multi-behavior view is paired with other behavior-specific views as positive pairs for the same user, while considering other users as negative samples.
  • Contrastive Objective. Most of the methods (Yang et al., 2021a; Gu et al., 2022; Wei et al., 2022; Li et al., 2023b; Qiao et al., 2023; Xuan and Li, 2023; Wei et al., 2023b; Xuan et al., 2023) in MbeRec utilize the InfoNCE objective to optimize contrastive learning. In MMCLR, it explicitly pulls the positive pairs closer together while pushing away negative pairs by ensuring that the similarity score of the former is larger than that of the latter.
4.8.3. Generative Learning in MbeRec
Generative learning in MbeRec (Table 8) adopts the variational autoencoding paradigm to reconstruct behavior interactions, thereby providing informative auxiliary learning signals.
  • Generative Learning Paradigm. Both VCGAE (Liu et al., 2023a) and BVAE (Rao et al., 2023) adopt the variational auto-encoding generative paradigm for modeling behavior heterogeneity. In VCGAE, a variation graph autoencoder is designed. It utilizes auxiliary behavior fusion to obtain the variance vector and a behavior transfer network to obtain the mean vector, which together form the distribution of latent factors. BVAE uses a behavior-aware semi-encoder to encode the variance vector based on users’ historical interactions. It learns the mean vector through a global feature filtering network, resulting in behavior-aware latent representations of users.
  • Generation Target. In the VAE paradigm, the latent representations obtained are fed into the decoder for generation. In VCGAE, the generation target is the edge in the user-item graph of the target behavior (e.g., purchase). Besides, BVAE focuses on generating the user interaction vector across different behaviors.
4.9. Multi-modal Recommendation
4.9.1. Task Formulation
In the context of Multi-modal Recommendation (MmoRec), auxiliary observed data $\mathcal{X}$ contains item multi-modal information. Typically, an item can possess auxiliary features such as text descriptions, images, or acoustic data. Each item is associated with modality-specific features $\mathbf{e}_{m}$, where $m\in\mathcal{M}$ represents modalities. MmoRec aims to utilize this multi-modal information to improve recommendation accuracy.
4.9.2. Contrastive Learning in MmoRec
Contrastive learning in MmoRec (Table 8) cleverly leverages the multi-modal information of items to construct different contrastive views. This enables effective fusion of multi-modal information and promotes the learning of recommenders.
  • View Creation. In data-based view creation methods, CMI (Li et al., 2022a) conducts data-level item sequence augmentation and encodes user interest embeddings to generate contrastive views. MMGCL (Yi et al., 2022) employs modality masking and edge dropout techniques to create augmented data views for each node in the multi-modal graph. Additionally, SLMRec (Tao et al., 2022b) utilizes data-level augmentation techniques such as feature dropout and masking to obtain two views for each node. In contrast, model-based methods like Liu et al. (Liu et al., 2022b) leverage a text/image encoder to encode two views for each item, which are then used for contrastive learning. Furthermore, MICRO (Zhang et al., 2022d) mines the latent graph structure of items and utilizes graph neural networks to encode modality-specific views and generate a fused view for each item. Meanwhile, MMSSL (Wei et al., 2023a) employs adversarial learning to obtain a modality-aware user-item graph and encodes modality-fused user embeddings as one view and another user’s view based on collaborative information. For feature-based view creation, BM3 (Zhou et al., 2023b) encodes modality embeddings, applies feature-level embedding dropout for augmented views, and includes an ID-based view from ID embeddings.
  • Pair Sampling. All the methods follow natural sampling. In (Liu et al., 2022b), positive pairs are formed by two views from different modalities of the same item, and negative pairs consist of views from different items. CMI creates positive pairs from two augmented interest-based views for the same user. MMGCL forms positive pairs from randomly augmented views and challenging negatives obtained by replacing modality data. SLMRec forms positive pairs with two data-augmented views of each node and ID-modality and other modality representations. MICRO maximizes agreement between item representations under individual modalities and fused multi-modal representations, forming positive pairs with modality-view and fused view of the same item and negative pairs with views from other items. BM3 aligns inter-modality by pairing augmented modality-views with the ID-based view and intra-modality by pairing modality-views of the same item. MMSSL forms positive pairs with the fused modality-view and collaborative-aware view of the same user, and negative pairs with views from other users.
  • Contrastive Objective. Most of the CL-based works utilize either InfoNCE-based (Li et al., 2022a; Yi et al., 2022; Tao et al., 2022b; Zhang et al., 2022d; Wei et al., 2023a) or JS-based (Liu et al., 2022b) loss functions as the loss function. Besides, BM3 directly aligns the views with cosine similarity.
4.9.3. Generative Learning in MmoRec
Generative learning, as recently shown by MVGAE (Yi and Chen, 2021) (Table 8), employs the graph VAE to transform modality data into a latent space and utilizes generation tasks to guide the learning process.
  • Generative Learning Paradigm. MVGAE utilizes a multi-modal variational graph auto-encoder to assign mean and variance vectors to each node in the user-item graph and employs the product-of-experts approach to fuse modality-specific latent representations for generation.
  • Generation Target. In MVGAE, the latent representations obtained from encoder are utilized to generate (reconstruct) the interaction edges in the user-item graph using an inner product decoder and BPR loss.
4.9.4. Adversarial Learning in MmoRec
In MmoRec, adversarial learning is recently used in MMSSL (Wei et al., 2023a) (Table 8) to perform structure learning based on multi-modal information, revealing modality-aware user-item interactions.
  • Adversarial Learning Paradigm. MMSSL generates a modality-aware user-item graph structure using a neural generator with item multi-modal features. Then, the user-specific interaction vector undergoes adversarial learning in a discriminator. The differentiable pipeline enables multi-modal adversarial user-item relation learning.
  • Discrimination Target. The user-item graph structure learned in MMSSL is split into user-specific interaction vectors, which are used for discrimination. The objective is to make these interaction vectors similar to real user interaction records in order to deceive the discriminator.
5. Discussions and Future Directions
In this section, we aim to delve into several open problems and potential future directions in the field of self-supervised learning for recommendation systems. By exploring and analyzing these challenges and opportunities, we aim to stimulate and encourage further research, development, and innovation in this rapidly advancing field of study.
5.1. Towards Foundation Recommender Models
Foundation model (Bommasani et al., 2021), trained on massive datasets and exhibiting remarkable generalization capabilities to handle a wide range of downstream tasks, has garnered significant attention from researchers across various domains (Radford et al., 2021; Bubeck et al., 2023; Xia et al., 2024; Tang et al., 2023). In recommender systems, current models using SSL techniques have shown significant performance, but they are limited by evaluation settings focused on a single dataset. Recent research in CV and NLP has improved models’ generalization abilities through self-supervised learning using contrastive learning (Radford et al., 2021) or generative learning (Touvron et al., 2023) on large amounts of data. A promising future direction is to design foundation recommender models that leverage SSL to learn user-item interaction patterns from massive data and achieve zero-shot cross-data reasoning and recommendation.
5.2. Unleashing the Potential of Denoised Diffusion Models
Denoised diffusion models have demonstrated remarkable proficiency in generating diverse types of data, including images, text, and even structured data. This new and popular generative learning paradigm has been effectively applied in various domains, as evidenced by recent studies such as (Rombach et al., 2022; Gong et al., 2022), and (Vignac et al., 2022). Recent studies in recommendation with generative self-supervised learning have started leveraging diffusion models either to generate augmented data (Liu et al., 2023b) or as the backbone model for inference (Wang et al., 2023d). We believe that the impressive generation capabilities demonstrated by diffusion models will bring new insights and fascinating models to the generative learning-enhanced recommendation in the near future, making it a promising research line.
5.3. Self-Supervised Integration of Large Language Models
Large language models (LLMs) have gained significant attention due to their exceptional performance in various domains (Zhao et al., 2023; Tang et al., 2024; Liu et al., 2024a). In recommender systems, LLMs can generate user/item profiles (Ren et al., 2024a; Wei et al., 2024a), explain user-item interactions (Li et al., 2023f), and serve as the backbone for recommendation with instruction tuning (Zhang et al., 2023a). The integration of large language models has brought about an unprecedented abundance of diverse, rich, and high-quality textual modality data. However, the challenge of effectively harnessing this textual capability remains an open research question. Recent works have employed self-supervised learning techniques, such as contrastive learning (Ren et al., 2024a) and mask-reconstruction (Wei et al., 2024a), to align the knowledge of LLMs with recommenders. As we look to the future, self-supervised learning techniques are poised to play a crucial role in enhancing recommenders with the incorporation of LLMs.
5.4. Self-Supervised Learning for Dynamic Recommendation Adaptation
Current recommendation research often assumes a fixed number of users and items, making it challenging for methods to adapt to continuous new data. In practice, models are deployed in dynamic environments with continuously generated user-item interactions and new users/items (You et al., 2022). To address this, prompt tuning (Lester et al., 2021; Wei et al., 2024b) is utilized to efficiently update pre-trained models on new data (Yang et al., 2024). Nevertheless, new data sparsity hinders effective supervision signals. Thus, effectively leveraging self-supervised learning for efficient learning on dynamic data is crucial.
5.5. Theoratical Foundation of various SSL paradigm
While self-supervised learning has improved recommender systems, a comprehensive theoretical foundation for various paradigms is lacking. Some methods have provided theoretical explanations for contrastive learning (Wu et al., 2021b; Ren et al., 2023b) and generative masked autoencoding in graphs (Li et al., 2022c). However, recent paradigms like denoised diffusion still need theoretical explanations to demonstrate their benefits in recommendation. This understanding would help uncover the underlying principles and mechanisms driving self-supervised learning in recommendation systems and offer insights into the generalization and robustness properties of these algorithms.
6. Conclusion